OpenStudy (anonymous):

2.What is the distance between the origin and (-3, -4)?

OpenStudy (anonymous):

\[ \sqrt{ x^2 + y^2} \]

OpenStudy (anonymous):

\[ \sqrt{ (-3)^2 + (-4)^2}=\sqrt {25}=5 \]